# Neural Eulerian Scene Flow Fields

### v2

@@ Abstract
We reframe scene flow as the task of estimating a continuous space-time ordinary differential equation (ODE) that describes motion for an entire observation sequence, represented with a neural prior. Our method, EulerFlow, optimizes this neural prior estimate against several multi-observation reconstruction objectives, enabling high quality scene flow estimation via self-supervision on real-world data. EulerFlow works out-of-the-box without tuning across multiple domains, including large-scale autonomous driving scenes and dynamic tabletop settings. Remarkably, EulerFlow produces high quality flow estimates on small, fast moving objects like birds and tennis balls, and exhibits emergent 3D point tracking behavior by solving its estimated ODE over long-time horizons. On the Argoverse 2 2024 Scene Flow Challenge, EulerFlow outperforms all prior art, surpassing the next-best unsupervised method by more than 2.5×2.5$\times$2.5 ×, and even exceeding the next-best supervised method by over 10%. See vedder.io/eulerflow for interactive visuals.
@@ 1 Introduction
Scene flow estimation is the task of describing motion with per-point 3D motion vectors between temporally successive point clouds (Dewan et al., 2016; Liu et al., 2019; Erçelik et al., 2022; Jund et al., 2021; Zhang et al., 2024b; Vedder et al., 2024; Khatri et al., 2024). Such per-point motion estimates are critical for autonomy in diverse environments, e.g., maneuvering around open-world objects like debris (Peri et al., 2022a) or folding deformable cloth (Weng et al., 2022). Importantly, scene flow estimation requires not only an understanding of object geometry, but also its motion. However, scene flow methods broadly do not work on smaller objects (Khatri et al., 2024). For example, in the autonomous vehicles domain, Khatri et al. highlight that even supervised methods struggle to describe the majority of pedestrian motion, with unsupervised methods failing dramatically. Scene flow promises to be a powerful primitive for understanding the dynamic world, but such failures explain why it has limited adoption in downstream applications like tracking (Zhai et al., 2020) or open-world object extraction (Najibi et al., 2022).
@@
Scene Flow via ODE. In Figure 2, visual assessment of scene flow quality is much easier in an accumulated global frame; while incomplete due to an implicit time axis, these accumulated flow vectors allow viewers to imagine how positions in 3D space evolve over many timesteps, and compare that to predicted flows. This imagination of scene flow as continuous motion over large time intervals motivates us to model scene flow as an ordinary differential equation (ODE) that describes the scene’s instantaneous motion across continuous position and time. Scene flow estimation then becomes the task of estimating this ODE. We can straightfowardly represent this ODE estimate with a neural prior (Li et al., 2021b) and optimize it against scene flow surrogate objectives, both over single frame pairs and extended across arbitrary time intervals to produce better quality estimates. We formalize this in Section 3 and propose the Scene Flow via ODE framework.
EulerFlow. We instantiate Scene Flow via ODE with standard point cloud distance objectives like Chamfer Distance to create EulerFlow. Notably, EulerFlow outperforms all prior art, supervised or unsupervised, on the Argoverse 2 2024 Scene Flow Challenge and Waymo Open Scene Flow benchmark. It outperforms prior unsupervised methods by a large margin (>2.5×>2.5$\times$> 2.5 × mean dynamic error reduction), and is able to capture small, fast moving objects, including those outside of labeled taxonomies (e.g. the flying bird in Figure 1(a)). Due to its simplicity, EulerFlow is able to provide high quality scene flow out-of-the-box on real-world data for other important domains such as dynamic tabletop settings (Figure 1(b)) without domain-specific tuning. Finally, we show that simple ODE solving techniques like Euler integration can be used to extract 3D point tracks (Figure 1(c)), which serves as both an exciting emergent behavior as well as a mechanism for visualizing and interpreting the quality of the continuous ODE estimate.
@@ 2 Background and Related Work
Evaluation. Dewan et al. formalized scene flow for point clouds as the task of estimating motion between point cloud $P_{t}{}$ at time $t$ and point cloud $P_{t+1}{}$ at $t+1$ by estimating the true flow $\mathcal{F}_{t,t+1}$, i.e. true residual vectors for every point in $P_{t}{}$ that describe its movement to its associated position at $t+1$. Error is computed by measuring the per-point endpoint distance between estimated and ground truth vectors. Historically, these errors are reported with a per-point average (Average EPE); however, as Chodosh et al. show, Average EPE is dominated by background points, preventing meaningful measurement of non-ego object motion descriptions. Khatri et al. address this shortcoming with Bucket Normalized EPE, which reports per-class performance normalized by speed, allowing for direct comparisons across classes with very different average speeds (e.g. pedestrians and cars). Bucket Normalized EPE is the basis for the Argoverse 2024 Scene Flow Challenge [^1] https://www.argoverse.org/sceneflow, where methods are ranked by the mean error of their motion descriptions (mean Dynamic Normalized EPE).
@@
Input / Output Formulation. Dewan et al.’s choice to formulate scene flow using only two input frames is arbitrary; it’s the minimal information needed to extract rigid motion, but there are not real-world problems constrained to only have access to two frames. Indeed, using five or ten frames of past observations is standard practice in the 3D detection literature (Zhu et al., 2019; Vedder & Eaton, 2022; Peri et al., 2022b; 2023; Nalty et al., 2022), and multi-frame formulations have started to appear in the scene flow literature: Liu et al. (2024) and Flow4D (Kim et al., 2024) use three ($P_{t-1},P_{t},P_{t+1}$) and five input frames ($P_{t-3},\ldots,P_{t+1}$) respectively to predict ^$\hat{\mathcal{F}}_{t,t+1}$. As we discuss in Section 3, rather than just using more observations to estimate flow for a single frame pair, we formulate scene flow as a joint estimation problem: given the full observation sequence $\left(P_{0},\ldots,P_{N}\right)$, we estimate all flows ^$\hat{\mathcal{F}}_{0,1},\ldots,\hat{\mathcal{F}}_{N-1,N}$ between all adjacent observations.
@@
Feedforward Methods. Feedforward networks are a popular class of scene flow methods due to their fast inference speed (Liu et al., 2019; Behl et al., 2019; Tishchenko et al., 2020; Kittenplon et al., 2021; Wu et al., 2020; Puy et al., 2020; Li et al., 2021a; Jund et al., 2021; Gu et al., 2019; Battrawy et al., 2022; Wang et al., 2022b; Kim et al., 2024; Zhang et al., 2024a). While they are often trained with supervised labels, recent work has developed distillation pipelines that leverage unsupervised pseudolabelers (Vedder et al., 2024; Zhang et al., 2024b; Lin & Caesar, 2024).
@@
Neural Scene Flow Prior. Li et al. (2021b) propose Neural Scene Flow Prior (NSFP), a widely adopted unsupervised scene flow approach. NSFP uses the inductive bias of the smooth, restricted learnable function class of two ReLU MLP coordinate networks (8 hidden layers of 128 neurons); $\theta$ to estimate forward flow and $\theta^{\prime}$ to estimate backwards flow, minimizing
[…]
@@
Motion Beyond Two Frames. Wang et al. (2022a) tackles the adjacent problem of estimating 3D point trajectories over 25 frames with Neural Trajectory Prior (NTP) by jointly optimizing three separate ReLU MLP neural priors: 1) a sinusoidal embedded, time conditioned, 25 frame trajectory basis estimator ($\textup{embed}(t)\mapsto 256\times 25\times 3$ tensor, where $256$ is the dimension of the trajectory basis), 2) a coordinate network bottleneck encoder, and 3) a bottleneck decoder to estimate a per-point linear combination over the learned trajectories. Trajectories are optimized for both a one-frame lookahead $L_{2}$ Chamfer loss and a cyclic consistency loss over the entire velocity space trajectory.
Deformation in Reconstruction. Nerfies (Park et al., 2021) and DynamicFusion (Newcombe et al., 2015) estimate a deformation field to warp a canonical frame to explain the observed frame. While capable of describing small motions, these methods require a canonical frame that contains all of the relevant geometry to deform; however, in large, highly dynamic scenes like autonomous driving, there is often no frame that contains all moving objects. By comparison, Scene Flow via ODE does not assume the existence of a canonical frame, instead only describing how the scene changes.
Figure 3: Overview of our Scene Flow via ODE framework, which estimates an ODE across the entire observation sequence by optimizing against multi-frame objectives. This ODE estimate is represented with a neural prior (Li et al., 2021b), providing a flexible, general representation for describing position-time motion.
@@ 3 Scene Flow via ODE
Formally, given a scene, let $L(x_{0},y_{0},z_{0},t)$ be the Lagrangian view of the scene’s true flow field, i.e. a continuous function that, based on a canonical frame at time 00, describes the true position of the canonical frame particle $x_{0},y_{0},z_{0}$ at some other time $t$. As we discuss in Section 2, this Lagrangian view is common in the the deformable reconstruction literature, and the requirement for a canonical frame definition means these approaches struggle to describe scenes where there is no frame that contains all moving objects.
@@ 4 EulerFlow
$\textup{Euler}_{\theta}\left(P_{t},k\right)=P_{t}+\mathcal{F}_{t,t+k}=\forall p\in P_{t}:E_{\theta}(p_{xt},p_{yt},p_{zt},t,t+k)\enspace,$ (2)
@@
$\arg\min_{\theta}\sum{\begin{array}[]{l}\forall k\in\{-W,\ldots,W\}\setminus\{0\}:\textup{TruncatedChamfer{}}(\textup{Euler}_{\theta}\left(P_{t},k\right),P_{t+k})\\ \alpha\left\lVert\textup{Euler}_{\theta}\left(\textup{Euler}_{\theta}\left(P_{t},1\right),-1\right)-P_{t}\right\rVert_{2}\end{array}}\enspace$ (3)
@@ 5 Experiments
In order to validate EulerFlow’s construction and better understand the impact of its design choices, we perform extensive experiments on the Argoverse 2 (Wilson et al., 2021) and Waymo Open (Sun et al., 2020) autonomous vehicle datasets. We compare against open source implementations of FastNSF (Li et al., 2023), Liu et al., NSFP (Li et al., 2021b), FastFlow3D (Jund et al., 2021), and variants of ZeroFlow (Vedder et al., 2024) provided by the ZeroFlow model zoo [^3] https://github.com/kylevedder/SceneFlowZoo, from Vedder et al. (2024)., a third-party implementation of NTP (Wang et al., 2022a) from Vidanapathirana et al., and Argoverse 2 2024 Scene Flow Challenge leaderboard submission results from the authors of Flow4D (Kim et al., 2024), TrackFlow (Khatri et al., 2024), DeFlow++/DeFlow (Zhang et al., 2024a), ICP Flow (Lin & Caesar, 2024), and SeFlow (Zhang et al., 2024b). As discussed in Khatri et al. and used in the Argoverse 2 2024 Scene Flow Challenge, methods are ranked by their speed normalized mean Dynamic Normalized EPE.
Implementation Details. To showcase the flexibility of EulerFlow without hyperparameter tuning, for all quantitative experiments we run with a neural prior of depth 8 (NSFP’s default depth), except for our submission to the Argoverse 2 2024 Scene Flow Challenge (Section 5.1) where, based on our depth ablation study on the val split (Section 5.2.3), we set the depth of the neural prior to 18. As discussed in NTP’s original paper (Wang et al., 2022a) and confirmed by our experiments, NTP struggles to converge beyond 25 frames, so we only compare against it in a 20 frame settings. As is typical in the scene flow literature (Chodosh et al., 2023), we perform ego compensation and ground point removal on both Argoverse 2 and Waymo Open using the dataset provided map and ego pose.
@@ 5.1 How does EulerFlow compare to prior art on real data?
EulerFlow achieves state-of-the-art performance on the Argoverse 2 2024 Scene Flow Challenge leaderboard. Despite being unsupervised, EulerFlow surpasses all prior art, supervised or unsupervised, including Flow4D (Kim et al., 2024) [^4] Flow4D is the winner of the 2024 Argoverse 2 Scene Flow Challenge supervised track. and ICP Flow (Lin & Caesar, 2024) [^5] ICP Flow is the winner of the 2024 Argoverse 2 Scene Flow Challenge unsupervised track.. Notably, EulerFlow achieves <2.5×<2.5$\times$< 2.5 × lower error mean Dynamic EPE than ICP Flow and beats Flow4D by over 10%.
@@
EulerFlow’s dominant performance also holds on Waymo Open (Sun et al., 2020); we compare against several popular methods (Figure 6), and EulerFlow again out-performs the baselines by a wide margin, more than halving the error over the next best method.
@@ 5.2.3 How does the capacity of the neural prior impact EulerFlow?
Li et al. ablate the capacity of NSFP’s neural prior to characterize underfitting and overfitting to optimization objective noise, ultimately selecting a depth of 8. EulerFlow’s neural prior is structured similarly; however, NSFP is fitting a single snapshot in time, while EulerFlow is fitting an entire ODE over significant time intervals. Intuitively, one would expect that full sequence modeling would benefit from greater network capacity.
@@
Figure 13: Visualization of EulerFlow compared to prior art for the same scene as Figure 1(a) and Figure 9(a). EulerFlow is able to extract the bird’s trajectory; however, all other methods except Liu et al. fail to recognize this motion, and Liu et al.’s flow is marred by severe scene artifacts. The bird is outside the labeled object taxonomy, and so its motion is unlabeled in the ground truth (Figure 13(f)).

### v3

@@ Abstract
We reframe scene flow as the task of estimating a continuous space-time ordinary differential equation (ODE) that describes motion for an entire observation sequence, represented with a neural prior. Our method, EulerFlow, optimizes this neural prior estimate against several multi-observation reconstruction objectives, enabling high quality scene flow estimation via self-supervision on real-world data. EulerFlow works out-of-the-box without tuning across multiple domains, including large-scale autonomous driving scenes and dynamic tabletop settings. Remarkably, EulerFlow produces high quality flow estimates on small, fast moving objects like birds and tennis balls, and exhibits emergent 3D point tracking behavior by solving its estimated ODE over long-time horizons. On the Argoverse 2 2024 Scene Flow Challenge, EulerFlow outperforms all prior art, surpassing the next-best unsupervised method by more than $2.5\times$, and even exceeding the next-best supervised method by over 10%. See vedder.io/eulerflow for interactive visuals.
@@ 1 Introduction
Scene flow estimation is the task of describing motion with per-point 3D motion vectors between temporally successive point clouds (dewan2016rigid; flownet3d; flowssl; scalablesceneflow; seflow; vedder2024zeroflow; khatri2024trackflow). Such per-point motion estimates are critical for autonomy in diverse environments, e.g., maneuvering around open-world objects like debris (peri2022towards) or folding deformable cloth (weng2022fabricflownet). Importantly, scene flow estimation requires not only an understanding of object geometry, but also its motion. However, scene flow methods broadly do not work on smaller objects (khatri2024trackflow). For example, in the autonomous vehicles domain, khatri2024trackflow highlight that even supervised methods struggle to describe the majority of pedestrian motion, with unsupervised methods failing dramatically. Scene flow promises to be a powerful primitive for understanding the dynamic world, but such failures explain why it has limited adoption in downstream applications like tracking (Zhai2020FlowMOT3M) or open-world object extraction (objectdetectionmotion).
@@
Scene Flow via ODE. In Figure 2, visual assessment of scene flow quality is much easier in an accumulated global frame; while incomplete due to an implicit time axis, these accumulated flow vectors allow viewers to imagine how positions in 3D space evolve over many timesteps, and compare that to predicted flows. This imagination of scene flow as continuous motion over large time intervals motivates us to model scene flow as an ordinary differential equation (ODE) that describes the scene’s instantaneous motion across continuous position and time. Scene flow estimation then becomes the task of estimating this ODE. We can straightfowardly represent this ODE estimate with a neural prior (nsfp) and optimize it against scene flow surrogate objectives, both over single frame pairs and extended across arbitrary time intervals to produce better quality estimates. We formalize this in Section 3 and propose the Scene Flow via ODE framework.
EulerFlow. We instantiate Scene Flow via ODE with standard point cloud distance objectives like Chamfer Distance to create EulerFlow. Notably, EulerFlow outperforms all prior art, supervised or unsupervised, on the Argoverse 2 2024 Scene Flow Challenge and Waymo Open Scene Flow benchmark. It outperforms prior unsupervised methods by a large margin ($2.5\times$ mean dynamic error reduction), and is able to capture small, fast moving objects, including those outside of labeled taxonomies (e.g. the flying bird in Figure 1(a)). Due to its simplicity, EulerFlow is able to provide high quality scene flow out-of-the-box on real-world data for other important domains such as dynamic tabletop settings (Figure 1(b)) without domain-specific tuning. Finally, we show that simple ODE solving techniques like Euler integration can be used to extract 3D point tracks (Figure 1(c)), which serves as both an exciting emergent behavior as well as a mechanism for visualizing and interpreting the quality of the continuous ODE estimate.
@@ 2 Background and Related Work
Evaluation. dewan2016rigid formalized scene flow for point clouds as the task of estimating motion between point cloud $P_{t}{}$ at time $t$ and point cloud $P_{t+1}{}$ at $t+1$ by estimating the true flow $\mathcal{F}_{t,t+1}$, i.e. true residual vectors for every point in $P_{t}{}$ that describe its movement to its associated position at $t+1$. Error is computed by measuring the per-point endpoint distance between estimated and ground truth vectors. Historically, these errors are reported with a per-point average (Average EPE); however, as chodosh2023 show, Average EPE is dominated by background points, preventing meaningful measurement of non-ego object motion descriptions. khatri2024trackflow address this shortcoming with Bucket Normalized EPE, which reports per-class performance normalized by speed, allowing for direct comparisons across classes with very different average speeds (e.g. pedestrians and cars). Bucket Normalized EPE is the basis for the Argoverse 2024 Scene Flow Challenge [^1] https://www.argoverse.org/sceneflow, where methods are ranked by the mean error of their motion descriptions (mean Dynamic Normalized EPE).
@@
Input / Output Formulation. dewan2016rigid’s choice to formulate scene flow using only two input frames is arbitrary; it’s the minimal information needed to extract rigid motion, but there are not real-world problems constrained to only have access to two frames. Indeed, using five or ten frames of past observations is standard practice in the 3D detection literature (cbgs; vedder2022sparse; peri2022futuredet; peri2023empirical; surveyofpersonrecog), and multi-frame formulations have started to appear in the scene flow literature: liu2024selfsupervisedmultiframeneuralscene and Flow4D (flow4d) use three ($P_{t-1},P_{t},P_{t+1}$) and five input frames ($P_{t-3},\ldots,P_{t+1}$) respectively to predict $\hat{\mathcal{F}}_{t,t+1}$. As we discuss in Section 3, rather than just using more observations to estimate flow for a single frame pair, we formulate scene flow as a joint estimation problem: given the full observation sequence $\left(P_{0},\ldots,P_{N}\right)$, we estimate all flows $\hat{\mathcal{F}}_{0,1},\ldots,\hat{\mathcal{F}}_{N-1,N}$ between all adjacent observations.
@@
Feedforward Methods. Feedforward networks are a popular class of scene flow methods due to their fast inference speed (flownet3d; behl2019pointflownet; tishchenko2020self; kittenplon2021flowstep3d; wu2020pointpwc; puy2020flot; li2021hcrf; scalablesceneflow; gu2019hplflownet; battrawy2022rms; 9856954; flow4d; zhang2024deflow). While they are often trained with supervised labels, recent work has developed distillation pipelines that leverage unsupervised pseudolabelers (vedder2024zeroflow; seflow; lin2024icp).
@@
Neural Scene Flow Prior. nsfp propose Neural Scene Flow Prior (NSFP), a widely adopted unsupervised scene flow approach. NSFP uses the inductive bias of the smooth, restricted learnable function class of two ReLU MLP coordinate networks (8 hidden layers of 128 neurons); $\theta$ to estimate forward flow and $\theta^{\prime}$ to estimate backwards flow, minimizing
$\textup{TruncatedChamfer{}}(P_{t}{}+\theta\left(P_{t}{}\right),P_{t+1})+\left\lVert P_{t}{}+\theta\left(P_{t}{}\right)+\theta^{\prime}\left(P_{t}{}+\theta\left(P_{t}{}\right)\right)-P_{t}\right\rVert_{2}\kern 5.0pt,$ (1)
@@
Motion Beyond Two Frames. ntp tackles the adjacent problem of estimating 3D point trajectories over 25 frames with Neural Trajectory Prior (NTP) by jointly optimizing three separate ReLU MLP neural priors: 1) a sinusoidal embedded, time conditioned, 25 frame trajectory basis estimator ($\textup{embed}(t)\mapsto 256\times 25\times 3$ tensor, where $256$ is the dimension of the trajectory basis), 2) a coordinate network bottleneck encoder, and 3) a bottleneck decoder to estimate a per-point linear combination over the learned trajectories. Trajectories are optimized for both a one-frame lookahead $L_{2}$ Chamfer loss and a cyclic consistency loss over the entire velocity space trajectory.
Deformation in Reconstruction. Nerfies (park2021nerfies) and DynamicFusion (dynamicfusion) estimate a deformation field to warp a canonical frame to explain the observed frame. While capable of describing small motions, these methods require a canonical frame that contains all of the relevant geometry to deform; however, in large, highly dynamic scenes like autonomous driving, there is often no frame that contains all moving objects. By comparison, Scene Flow via ODE does not assume the existence of a canonical frame, instead only describing how the scene changes.
Figure 3: Overview of our Scene Flow via ODE framework, which estimates an ODE across the entire observation sequence by optimizing against multi-frame objectives. This ODE estimate is represented with a neural prior (nsfp), providing a flexible, general representation for describing position-time motion.
@@ 3 Scene Flow via ODE
Formally, given a scene, let $L(x_{0},y_{0},z_{0},t)$ be the Lagrangian view of the scene’s true flow field, i.e. a continuous function that, based on a canonical frame at time 0, describes the true position of the canonical frame particle $x_{0},y_{0},z_{0}$ at some other time $t$. As we discuss in Section 2, this Lagrangian view is common in the the deformable reconstruction literature, and the requirement for a canonical frame definition means these approaches struggle to describe scenes where there is no frame that contains all moving objects.
@@ 4 EulerFlow
$\textup{Euler}_{\theta}\left(P_{t},k\right)=P_{t}+\mathcal{F}_{t,t+k}=\forall p\in P_{t}:E_{\theta}(p_{xt},p_{yt},p_{zt},t,t+k)\kern 5.0pt,$ (2)
@@
$\arg\min_{\theta}\sum{\begin{array}[]{l}\forall k\in\{-W,\ldots,W\}\setminus\{0\}:\textup{TruncatedChamfer{}}(\textup{Euler}_{\theta}\left(P_{t},k\right),P_{t+k})\\ \alpha\left\lVert\textup{Euler}_{\theta}\left(\textup{Euler}_{\theta}\left(P_{t},1\right),-1\right)-P_{t}\right\rVert_{2}\end{array}}\kern 5.0pt$ (3)
@@ 5 Experiments
In order to validate EulerFlow’s construction and better understand the impact of its design choices, we perform extensive experiments on the Argoverse 2 (argoverse2) and Waymo Open (waymoopen) autonomous vehicle datasets. We compare against open source implementations of FastNSF (fastnsf), liu2024selfsupervisedmultiframeneuralscene, NSFP (nsfp), FastFlow3D (scalablesceneflow), and variants of ZeroFlow (vedder2024zeroflow) provided by the ZeroFlow model zoo [^3] https://github.com/kylevedder/SceneFlowZoo, from vedder2024zeroflow., a third-party implementation of NTP (ntp) from vidanapathirana2023mbnsf, and Argoverse 2 2024 Scene Flow Challenge leaderboard submission results from the authors of Flow4D (flow4d), TrackFlow (khatri2024trackflow), DeFlow++/DeFlow (zhang2024deflow), ICP Flow (lin2024icp), and SeFlow (seflow). As discussed in khatri2024trackflow and used in the Argoverse 2 2024 Scene Flow Challenge, methods are ranked by their speed normalized mean Dynamic Normalized EPE.
Implementation Details. To showcase the flexibility of EulerFlow without hyperparameter tuning, for all quantitative experiments we run with a neural prior of depth 8 (NSFP’s default depth), except for our submission to the Argoverse 2 2024 Scene Flow Challenge (Section 5.1) where, based on our depth ablation study on the val split (Section 5.2.3), we set the depth of the neural prior to 18. As discussed in NTP’s original paper (ntp) and confirmed by our experiments, NTP struggles to converge beyond 25 frames, so we only compare against it in a 20 frame settings. As is typical in the scene flow literature (chodosh2023), we perform ego compensation and ground point removal on both Argoverse 2 and Waymo Open using the dataset provided map and ego pose.
@@ 5.1 How does EulerFlow compare to prior art on real data?
EulerFlow achieves state-of-the-art performance on the Argoverse 2 2024 Scene Flow Challenge leaderboard. Despite being unsupervised, EulerFlow surpasses all prior art, supervised or unsupervised, including Flow4D (flow4d) [^4] Flow4D is the winner of the 2024 Argoverse 2 Scene Flow Challenge supervised track. and ICP Flow (lin2024icp) [^5] ICP Flow is the winner of the 2024 Argoverse 2 Scene Flow Challenge unsupervised track.. Notably, EulerFlow achieves $2.5\times$ lower error mean Dynamic EPE than ICP Flow and beats Flow4D by over 10%.
@@
EulerFlow’s dominant performance also holds on Waymo Open (waymoopen); we compare against several popular methods (Figure 6), and EulerFlow again out-performs the baselines by a wide margin, more than halving the error over the next best method.
@@ 5.2.3 How does the capacity of the neural prior impact EulerFlow?
nsfp ablate the capacity of NSFP’s neural prior to characterize underfitting and overfitting to optimization objective noise, ultimately selecting a depth of 8. EulerFlow’s neural prior is structured similarly; however, NSFP is fitting a single snapshot in time, while EulerFlow is fitting an entire ODE over significant time intervals. Intuitively, one would expect that full sequence modeling would benefit from greater network capacity.
@@
Figure 13: Visualization of EulerFlow compared to prior art for the same scene as Figure 1(a) and Figure 9(a). EulerFlow is able to extract the bird’s trajectory; however, all other methods except liu2024selfsupervisedmultiframeneuralscene fail to recognize this motion, and liu2024selfsupervisedmultiframeneuralscene’s flow is marred by severe scene artifacts. The bird is outside the labeled object taxonomy, and so its motion is unlabeled in the ground truth (Figure 13(f)).